### v1

@@ 6.4 Functional equations
For $10<\lambda<$ 1, the theorem of Koenigs (1884) states that there is a unique solution holomorphic in a neighbourhood of the origin satisfying $u(0)=0,u^{\prime}(0)=1$. Here will will approximate this solution on a finite interval by interpolating at scaled Chebyshev points and employing the barycentric resampling matrix as described in previous sections.
Figure 17 shows code for implementing the above on the domain $[0,\pi]$ with $\lambda=\tfrac{1}{2}$ and compares the result with the iterated composition solution $u(t)=\lim_{n\rightarrow\infty}\lambda^{-n}f^{\circ n}(t)$, where $f^{\circ n}(t):=\underbrace{f(f(f(\ldots)))}_{n\text{ times}}.$ As with the FDE examples of previous sections, we observe geometric convergence as the number of interpolation points is increased. The interpolation approach has the advantage that it may be applied to more general FEs where iterated solutions are not readily available.
@@
3P = barymat(f(t), t, w); I = eye(n)
4L = P - lam*I;
5[~, ~, V] = svd(L); u = V(:,end);
6D = diffmat(t, w);  u = u./(D(1,:)*u);
Refer to captionRefer to caption
Figure 17: Left: MATLAB code to solve (6.4) using the barycentric resampling appraoch. The final line normalises to ensure that $u^{\prime}(0)=0$. Centre: Agreement of the computed solutions using interpolation and iterated composition. Right: Geometric convergence of the interpolation solution as the discretisation size is increased.
@@ 7 Conclusion
The barycentric resampling matrix was used to create a framework enabling the solution of various delay and functional differential equations—including pantograph, discrete, continuous, and state-dependent delay—via Chebyshev spectral collocation. Furthermore, this framework was shown to be application neutral-type equations, higher-order DDEs/FDEs, and DDE eigenvalue problems. The close connection to standard Chebyshev spectral methods for ODEs enables the ready implementation of the method to existing code bases, as was demonstrated here in the case of Chebfun/Chebop. Numerous examples were presented, demonstrating both rapid (typically exponential) convergence as well as convenient implementation. Furthermore, the method was shown to be readily extendable to methods based on other classical orthogonal polynomials, trigonometric polynomials, or rational functions. We close with a few additional observations and comments.
In the coded examples given, boundary bordering was used to implement boundary conditions. If rectangular projection is used, a la (Driscoll and Hale, 2016), then one could explicitly construct the discrete evaluation operation operators to return values on the projected grid, thus avoiding an addition matrix multiplication. Formulas for the explicit construction of the rectangular differentiation matrices are also known (Xu and Hale, 2016).
@@
Here we employed spectral collocation, which typically gives rise to dense linear systems, requiring $O(n^{3})$ floating point operations to solve an $n$-point discretisation using dense linear algebra. By utilising the connection to discrete cosine transform (DCT) and the FFT, Chebyshev differentiation matrices can be applied in $\mathcal{O}(n\log(n))$ operations, making iterative methods like MINRES/GMRES (with suitable preconditioners) attractive (Canuto and Quarteroni, 1985). In general, the interpolation operator/barycentric resampling matrix requires $O(n^{2})$ operations to apply directly, making an iterative solver less attractive in the DDE/FDE setting. However, this can be mitigated by the use of a non-uniform DCT (Dutt and Rokhlin, 1993; Potts and Steidl, 2003).
@@
An alternative to spectral collocation methods are spectral methods based on coefficient expansions (e.g., Galerkin or tau methods). The ultraspherical spectral method is such an example, which gives rise to sparse (almost-banded) matrices when applied to the ODEs with polynomial coefficients (Olver and Townsend, 2013). One could introduce an interpolation operator to these discretisations, similarly to that in Section 3. This would destroy the sparseness of the discretisation, but other beneficial properties, such as improved conditioning, would likely be maintained, making such an approach attractive. We leave further investigations for future work.
@@ Acknowledgements
The author is grateful to André Weideman and Toby Driscoll for useful discussions.

### v2

@@ 6.4 Functional equations
For $10<\lambda<$ 1, the theorem of Koenigs (1884) states that there is a unique solution holomorphic in a neighbourhood of the origin satisfying $u(0)=0,u^{\prime}(0)=1$. Here we approximate this solution on a finite interval by interpolating at scaled Chebyshev points and employing the barycentric resampling matrix as described in previous sections, yielding a system $\big{(}P(f(\bm{t});\bm{t})-\lambda I\big{)}\bm{u}=0$. This can be augmented with the boundary constraint, $D_{1,:}\bm{u}=1$, and then solved for $\bm{u}\approx u(\bm{t})$. Figure 16 shows code for implementing the above on the domain $[0,\pi]$ with $\lambda=\tfrac{1}{2}$ and compares the result with the iterated composition solution $u(t)=\lim_{n\rightarrow\infty}\lambda^{-n}f^{\circ n}(t)$, where $f^{\circ n}(t):=\underbrace{f(f(f(\ldots)))}_{n\text{ times}}.$ As with the FDE examples of previous sections, we observe geometric convergence as the number of interpolation points is increased. The interpolation approach has the advantage that it may be applied to more general FEs where iterated solutions are not readily available.
@@
3P = Barymat(f(t), t, v); I = eye(n)
4A = [D(1,:) ;
5 P(2:end,:) - lam*I(2:end,:);
6rhs = [1 ; zeros(n-1,1)];
[…]
Figure 16: Left: MATLAB code to solve the functional equation (6.7) using the barycentric resampling approach. Centre: Agreement of the computed solutions using interpolation and iterated composition. Right: Geometric convergence of the interpolation solution to the iterated solution as the discretisation size is increased.
@@ 7 Conclusion
The barycentric resampling matrix was used to create a framework enabling the solution of various delay and functional differential equations—including pantograph, discrete, continuous, and state-dependent delay—via Chebyshev spectral collocation. Furthermore, this framework was shown to be applicable to neutral-type equations, higher-order DDEs/FDEs, and DDE-EVPs. The close connection to standard Chebyshev spectral methods for ODEs enables the ready implementation of the method to existing code bases, as was demonstrated here in the case of Chebfun/Chebop. Numerous examples were presented, demonstrating both rapid (typically geometric) convergence as well as convenient implementation. Furthermore, the method was shown to be readily extendable to methods based on other classical orthogonal polynomials, trigonometric polynomials, or rational functions. We close with a few additional observations and comments.
For IVP-DDES, it is unlikely that the proposed method will competitive with standard time-stepping–plus–interpolation techniques, for the same reason that spectral methods are rarely used for IVP-ODES (i.e., waveform relaxation methods). However, the example in Section 5.2 demonstrates that there may still be some stability benefits to such a global approach for some DDE-IVP problems. For DDE-BVPS, PDDEs, DDE-EVPs, or FDEs that cannot be solved with a method of steps-type approach, there are likely benefits to a spectral collocation method in terms of efficiency and/or accuracy when compared to other approaches—be they shooting-type methods or lower-order collocation approaches—particularly when a high degree of accuracy is required.
In the examples above, we dedicated little discussion to the computational considerations of solving the linear systems of equations resulting from the spectral collocation. For the most part, this is because our focus was, in spectral element terminology, on the "high $p$, large $h$" regime (i.e., high degree polynomials on a coarse mesh), giving rise to dense matrices and $\mathcal{O}(n^{3})$ solves when using direct methods. However, there is perhaps still structure than can be exploited in some cases, particularly as the number of subintervals is increased. Here we discuss some such possibilities.
In Example 4 we mentioned that the resulting block lower triangular (BLT) discretisation could be efficiently solved via the method of steps for an IVP-DDE. For a BVP-DDE, the augmented boundary conditions will typically result in an ‘almost block lower triangular’ system (i.e., BLT with small number of dense rows). The BLT structure can then be exploited by an appropriate Schur-complement factorisation or the Woodbury formula. These ideas may also be applicable in the FDE case when $\tau(t)\approx t$, leading to almost block-banded systems that could be efficiently solved in a similar way. Another situation that might result in better efficiency for some autonomous problems with periodic boundary conditions, discretised on uniform intervals with the same discretisation size. In such cases, the resulting linear system has a block circulant structure, which can be exploited (De Mazancourt and Gerlic, 1983). However, a global trigonometric interpolant, as in Example 11, may be better suited to such cases.
@@
Rather than direct methods for solving the linear systems, we might consider iterative approaches. In particular, by utilising the connection to the discrete cosine transform (DCT), Chebyshev differentiation matrices can be applied in $\mathcal{O}(n\log(n))$ operations, making methods like BiCG or GMRES (with suitable preconditioners) attractive (Canuto and Quarteroni, 1985). In general, the interpolation operator/barycentric resampling matrix requires $\mathcal{O}(n^{2})$ operations to apply directly, making an iterative solver less attractive in the DDE/FDE setting. However, this can perhaps be mitigated by the use of a non-uniform DCT (Potts and Steidl, 2003) or fast multipole methods (Dutt et al., 1996).
In the coded examples given, boundary bordering was used to implement boundary conditions. If rectangular projection is used, a la (Driscoll and Hale, 2016), then one could explicitly construct the discrete evaluation operators to return values on the projected grid, thus avoiding an additional matrix multiplication. Formulas for the explicit construction of the rectangular differentiation matrices are also known (Xu and Hale, 2016).
@@
An alternative to spectral collocation methods are spectral methods based on coefficient expansions (e.g., Galerkin or tau methods). The ultraspherical spectral method is such an example, which gives rise to sparse (almost-banded) matrices when applied to ODEs with polynomial coefficients (Olver and Townsend, 2013). One could introduce an interpolation operator to these discretisations, similarly to that in Section 3. This would destroy the sparseness of the discretisation, but other beneficial properties, such as improved conditioning, would likely be maintained, making such an approach attractive.
Finally, rather than the “high $p$, large $h$” approach, one could instead interpolate with lower degree interpolants on a finer mesh, trading exponential convergence with dense matrices for algebraic convergence with sparse matrices. The latter is essentially the approach taken by DDE-BIFTOOL (Sieber et al., 2016). A detailed comparison of the two approaches is beyond the scope of this paper, but the most efficient approach will likely depend on both the accuracy requirements and on the problem at hand. For smooth problems, such as Example 9 or PDDEs with smooth coefficients, it is typically difficult to beat the geometric convergence of spectral methods by introducing artificial discontinuities in the discretisation (Trefethen, 2000, p. x). For non-smooth problems with multiple discontinuities present in the solution itself, for example, a non-periodic discrete delay DDE-BVP with $p\ll T$, a lower-order discretisation will likely be more efficient.
Example 13.
To demonstrate the former case, we close with one final example, namely computing a limit cycle of the delayed Logistic equation
$y^{\prime}(t)=(\lambda-y(t-1))u(t),\qquad y(t+T)=y(t),\quad\forall t.$ (7.1)
This example is considered in (Engelborghs et al., 2000a), where a piecewise 3-point Gauss–Legendre discretisation was used to calculate a stable limit cycle with period $T=4.0964$ for $\lambda=1.7$. Using $200$ uniform subintervals, an accuracy of around $5\times 10^{-9}$ throughout the interval of integration is achieved, equating to 600 degrees of freedom. As shown in Figure 17 below, the same accuracy can be achieved using around 50 degrees of freedom with a global polynomial interpolant and only 25 degrees of freedom with a global trigonometric polynomial interpolant. [^11] In (Engelborghs et al., 2000a) the error at the interpolation nodes is shown to be lower, about $10^{-11}$, due to super convergence. However, the number of degrees of freedom required to achieve this accuracy in the spectral methods is only increased marginally, due to the geometric convergence. Of course, degrees of freedom are not the whole story, as one may more readily take advantage of sparsity and or adaptivity in the piecewise discretisation. However, this example suggests that there may be some benefit to global discretisations for such problems. We hope to pursue further investigations in future work.
Figure 17: Left: Periodic solution computed using the spectral collocation method. An initial approximation to the periodic solution is obtained using dde23, with the initial guess $T=4$ and assuming $y(t\leq 0)=0.5$. This is then refined using the approach described in Example 11. Centre: Computed limit cycle in the $(y,y^{\prime})$ plane (solid) and the initial trajectory from dde23 (dashed). Right: Geometric convergence of the spectral collocation approach using both Chebyshev and trigonometric interpolants. (The $n=70$ and $n=40$ solutions are respectively used as a proxy for the unknown analytic solution.)
@@ Acknowledgements
The author is grateful to André Weideman and Toby Driscoll for useful discussions and to the anonymous referees for their detailed and constructive feedback.